5. A jar contains 24 marbles, some are green and others are blue. If a marble is drawn at

random from the jar, the probability that it is green is ⅔. Find the number of blue balls

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Total marbles = 24 Let the total green marbles = x So, the total blue marbles = 24-x P(getting green marble) = x/24 From the question, x/24 = ⅔ So, the total green marbles = 16 And, the total blue marbles = 24-x = 8